OpenStudy (anonymous):

Solve for s -30 = 3(s + -34)

OpenStudy (paki):

divide both sides of the equation by "3" and show ur work...

OpenStudy (anonymous):

s=24

OpenStudy (paki):

how...? can u show ur work here?

OpenStudy (anonymous):

expand then follow pemdas. distibute parenthesis =3s-3*34 refine =3s-102 3s-102=-30 add 102 to both sides 3s-102+102=-30+102 3s=72 divide both sides by 3 3s/3 72/3 s=24

OpenStudy (paki):

yeah correct... :)